OpenStudy (anonymous):

solve the enequality 9(x-3)<4(x+10)

OpenStudy (mathmale):

Using the distributive property of multiplication, multiply out 9(x-3) and 4(x+10). Show all of your work, please.

OpenStudy (anonymous):

here are some choices: x<13/5 x>13/5 x<67/5 x>67/5

OpenStudy (mathmale):

123: Please follow my advice. Multiply out those two products.

OpenStudy (anonymous):

i dont know how to multiply

OpenStudy (anonymous):

well i kinda do

OpenStudy (unklerhaukus):

for the lefthand side 9(x-3) = 9*x +9*-3 =

OpenStudy (mathmale):

Multiply x by 9. We don't know the numerical value yet. Write out your answer as an algebraic expression. See Unkle's demo, above.

OpenStudy (anonymous):

i dont understand

OpenStudy (unklerhaukus):

what do you get when you simplify the left hand side?

OpenStudy (anonymous):

9 -3x

OpenStudy (mathmale):

2*3=6 is an example of multiplication of two numerals. 2*a = 2a is an exaple of multiplication of a letter (a quantity) by a numeral. 9(x)=?

OpenStudy (mathmale):

 ... Now please multiply out 4(x+10).